1) 2 to the power x=1-x² solve for x. 2)|x²+x+1|=3 where x€R ,find x

Grade:12

1 Answers

Arun
6 years ago
2)-
|x2 + x+ 1| = 3
now when we open mod, there are two cases-
1st Case-
x2 + x+ 1 = 3
x2 + x –2 = 0
x =  – 2,  x = 1
2nd Case-
x2 + x+ 1 = –3
x2 + x+ 4 = 0
in this equation D
hence roots are imaginary.